### v1

@@ V-F Summary
Given the privacy, computational and cost constraints, multiple LLMAs need to interact with each other to perform sequential Bayesian inference on online platforms. This section studied Bayesian social learning in a sequence of LLMAs to analyze multiple LLMAs performing sequential state estimation on online platforms. We discuss the relaxation of the standard Bayesian social learning protocol, in which the agents are allowed to share their private observations. We illustrate the effect on the herding threshold of the public prior, when LLMAs are allowed to share private observations and use more accurate LLMs as sensors.
@@ VI Asymmetric Information Structures of Large Language Model Agents
Motivated by the observed model collapse while training LLMs in a sequential manner and data incest in asynchronous Bayesian sensors, this section studies word-of-mouth and asynchronous social learning. Word-of-mouth social learning is a hierarchical social learning paradigm characterized by asymmetric information flow, where lower-level agents process and communicate observations, and top-down influence, where top-level agents dictate the public prior. We provide protocols for both kinds of social learning and a corollary that shows that information cascades happen in word-of-mouth learning as well. Techniques to perform stochastic control to prevent model collapse and data incest are not in the scope of this paper and can be considered in future work.
@@
There has been a lot of research recently that shows that LLMs exhibit model collapse when iteratively trained on data generated by another LLM [67]. In model collapse, the output probability distribution of the model collapses to a degenerate distribution as the model is trained iteratively on data generated by the previously trained model [37]. Similar results have been shown on model distillation, where a smaller LLM is trained using data generated from a larger LLM [79]. We use the Bayesian social learning model to show how an information cascade in LLMAs is similar to the model collapse observed while training LLMs.
@@
Asynchronous social learning in LLMAs is motivated by real-time settings like online platforms where there is a stream of data of the order of a hundred thousand every second [52, 4, 17]. Since the LLMs of LLMAs often take milliseconds to process a query, especially if these are third-party services, sequential Bayesian learning is often not possible. This is true particularly when the LLMAs are used for tasks that are more sophisticated than just Bayesian inference[49].
@@ VI-B Word of Mouth Bayesian Social Learning in LLM Agents
We now describe the word-of-mouth social learning protocols in $M$ LLMAs. The protocol is summarized as a pseudo-code in Algorithm 5. The protocol can be considered to run on two timescales. On the slower time scale, the first LLMA receives a new text observation $z_{kM}$ of the state $x$. On the slower time scale, the LLMAs communicate with each other in a sequential fashion by generating text observations corresponding to the low-dimensional features from the received text observation. That is, agent $m$ first receives a text observation $z_{m}$ and uses an LLM to obtain a low dimensional observation $y_{m}$. The agent $m$ then takes an action to update the public prior using (14). Then, the LLMA uses the LLM to generate a synthetic text observation $z_{m+1}$ used by the next agent.
@@
2:     LLMA-($kM$) receives a observation $z_{kM}\sim\mathds{P}(z|x)$.
3:     for $m\in kM,kM+1,kM+2,\dots kM-2$ do
@@
The above result can be proved directly using Theorem 2, where each agent has a different observation likelihood based on the previous agent, each of which is a concatenation of two observation likelihoods: the LLM as a low-dimensional sensor map and the LLM sampling a text observation from the text observation. Corollary 1 shows that even in the modified protocol of Algorithm 5, where synthetic data is used to aid the decision-making of LLMAs, cascades are inevitable.
@@
Asynchronous Social Learning in LLM Agents: We finally consider the asynchronous social learning setting in LLMAs. Here, the main difference between the previous two topologies is that the agents don’t necessarily act in a predefined sequential manner and don’t necessarily coordinate.
The protocol is summarized as a pseudo-code in Algorithm 6. The public prior $\pi_{k}$ is updated asynchronously when a LLMA receives a new private observation $z_{k}$, which it parses using the LLM to obtain a low-dimensional observation $y_{k}$. This observation is used to compute the posterior at time $\pi_{k}$ using a previous prior $\pi_{k-1}$. The posterior is then broadcasted.
@@ VII Summary
Part III, therefore, looks at stochastic control for LLMAs. For both regimes of LLMAs, when they are collaborative and autonomous, the paper formulates optimal stopping time problems to control herding by balancing the tradeoff between privacy and estimation. Structural assumptions on the optimal policy of the stopping time problems are proved by making structural assumptions on the cost and observation probabilities. The proposed solutions are extensions to our work in quickest change detection and quickest time herding [39, 41, 30]. A stochastic approximation algorithm is proposed to estimate the optimal policy without the knowledge of the transition probabilities.
@@ VIII Optimal Stopping Time Control in Centrally Controlled LLMAs
When LLMAs are deployed in real-life settings, they often exhibit bias in their actions, especially when there are multiple such agents [9]. The previous section showed that such bias could be explained by the herding behavior of LLMAs. This section formulates an optimal stopping time problem for quickest time herding in a sequence of LLMAs, which ensures that the herding is optimally delayed by letting the LLMAs share private observations. The delay in herding helps improve the state estimate. We discuss a stochastic control approach to solve the optimal stopping time problem and state assumptions that ensure that the optimal policy for the stopping time problem has a switching threshold curve with respect to the public prior. This structural result is exploited in Section X to efficiently approximate the optimal policy. Finally, we discuss extensions of the problem framework to the optimal switching between models of different sizes in sequential state estimation tasks is discussed. The schematic of the system model considered is illustrated in Figure 10.
@@
VIII-B Socialistic Objective For Optimal Stopping
@@
This subsection formulates an optimal stopping time problem to delay herding by making the agents optimally switch between two modes, acting benevolently by sharing their private observations or herding by performing the action from (14). Let $a_{k}\in\{0=\texttt{share private observation},1=\texttt{herd}\}$ denote the decision at $k$ for the chosen mode. Let $\mu:\mathcal{P}(\mathcal{X})\to\{0=\texttt{share private observation},1=\texttt{herd}\}$ denotes the stationary policy which maps the public prior to the decision rule. $\mu$ is a sufficient statistic for optimally delaying herding since the information cascade depends on only the public prior [40]. For this section, we assume that $\mathcal{X}=\mathcal{U}=\mathcal{Y}$, that is, in the sequential detection task, the observation $y$ is the noisy observation of the underlying state $x$ and $u$ is the maximum a-priori estimate of the state. Further, we formulate the stopping time problem such that one of the states is of special interest (denoted by $e_{1}$), and a decision to herd is taken once this state is estimated.
@@
To formulate the objective for the optimal stopping time problem, we consider the following natural filtration, $\mathcal{F}_{k}$ of the actions and decisions till time $k$,
$\displaystyle\mathcal{F}_{k}=\sigma(\{u_{1},\dots,u_{k-1},a_{1},\dots,a_{k}\}).$
@@
The optimal stopping time problem is to decide when to stop sharing private observations ($a_{k}=0$) and announce state $e_{1}$ ($a_{k}=1$). Let $\tau$ denote the stopping time with respect to the filtrations $\mathcal{F}_{k},k=1,\dots$. We now state the socialistic objective that each LLMA optimizes to solve the optimal stopping time problem and achieve quickest time herding,
$\displaystyle\begin{split}J(\mu)=\mathds{E}_{\mu}&\left\{\sum_{k=1}^{\tau-1}\rho^{k-1}\mathds{E}\left\{c(x,u_{k})|\mathcal{F}_{k-1}\right\}\right.\\ &\left.+\sum_{k=1}^{\tau-1}\rho^{k-1}d\mathds{E}_{\mu}\left\{\mathds{1}(x=e_{1})|\mathcal{F}_{k-1}\right\}\right.\\ &\left.+\ \rho^{\tau-1}\Upsilon\mathds{E}_{\mu}\left\{\mathds{1}(x\neq e_{1})\right\}\right.\\ &\left.+\ \frac{\rho^{\tau-1}}{1-\rho}\min_{u\in\mathcal{U}}\mathds{E}\left\{c(x,u)|\mathcal{F}_{\tau-1}\right\}\right\},\end{split}$ (19)
@@
Justification for socialistic cost: In the above expression, the first part corresponds to the discounted cost incurred if the first $\tau$ agents perform sensing, and the last term corresponds to the cost incurred for the agents after the stopping time $\tau$ (which herd and take the same opportunistic action). The second term is the delay cost with delay parameter $d$ in announcing if the underlying state is $e_{1}$ and the third term is the error cost for misclassification with $\Upsilon$ as the parameter. The socialistic cost incentivizes revealing private observations to delay herding so that enough private observations are available for estimation of the state. The reason we consider estimation with respect to a single state is that many practical applications focus on the identification of a critical state ( hateful user, bad product, etc.). The same socialistic cost function of (19) is optimized to obtain a policy with respect to the belief which decides when to stop herding.
@@
We consider the decision rule such that depending on the prior, the learners either share the private observation $y_{k}$ or they output the action which minimizes the expected cost of (14) under the public prior, i.e., the action $u_{k}$ is of the form,
@@
The motivation for this decision rule is two-fold: a) the herding can be delayed by ensuring enough private observations are shared, and b) the agents can reduce cost by outputting the opportunistic action once public prior is strong enough.
@@ VIII-C Structural Results on Optimal Policy
We now state the main structural result on the threshold structure of the optimal policy ((21)).
@@
Consider the sequential decision problem of LLMAs for detecting state $e_{1}$ with the social welfare cost of (19) and the constrained decision rule of (20). Then, under Assumption (S1-4), the constrained decision rule of (20) is a threshold in the belief space with respect to a threshold switching curve that partitions the belief space $\mathcal{P}(X)$. The optimal policy can be given by,
@@
The above theorem proves that under certain conditions on the cost function and observation matrix (S1-4), the optimal policy for solving the discounted socialistic cost optimal stopping time problem has a switching threshold curve. Such a switching threshold can be approximated by set of lines which can be searched for efficiently using a stochastic approximation based approach [40]. This is computationally tractable, in contrast to optimizing for a non-structured policy since it’s a finite approximation of a general infinite-dimensional policy.
@@ VIII-D Optimal Switching Between Different LLMs
The formulation that we proposed in this section can also be used by the LLMAs to switch between LLMs of different sizes. LLMs with a higher number of parameters can process more tokens and give more accurate responses; however, they are more expensive and use more computational resources. Therefore we can use an action space $\mathcal{U}=\{1=\mathcal{L}_{\text{small}},\mathcal{L}_{\text{large}}\}$, where $\mathcal{L}_{\text{large}}$ is an LLM with significantly (10x) more parameters than $\mathcal{L}_{\text{small}}$.
@@ VIII-E Summary
When a central entity is designing and deploying LLMAs for sequential Bayesian inference on online platforms, the entity needs to ensure such LLMAs don’t have a bias due to herding. In this section, we solved the problem of LLMAs optimally herding in Bayesian social learning framework to announce a particular state, such that their opportunistic cost and cost of obtaining and sharing the observation is balanced. The Bayesian agents considered in this section were cooperative and shared the same socialistically optimal policy. We also briefly discussed how such a scheme can be used to optimally switch between different sizes of LLMs to achieve an optimal tradeoff between cost and accuracy of estimation. In the next section we consider the problem in a different setting, where a central controller can incentivize autonomous LLMAs, and needs to control the incentives for improved state estimation.
@@ IX Optimal Stopping Time Control for Autonomous LLMAs
The LLMAs used for Bayesian inference can often be from different third-party services, each of which requires an incentive to perform the task. Motivated by controlling bias in such LLMAs, this section considers the problem of the central learner optimally incentivizing a sequence of autonomous LLMAs to delay their herding and obtain more accurate estimates. We first formulate the optimization problem of the central controller as a discounted cost POMDP and then show that under structural assumptions on the cost and observation matrix, the optimal policy has a threshold structure with respect to the public prior. This structure is exploited in Algorithm 7 to approximate the optimal incentivization policy of the central controller. The schematic of the setup considered in this section is illustrated in Figure 11.
@@
Figure 11: Schematic of stochastic control approach for optimally delaying herding in a sequence of incentivized autonomous LLM agents. This is motivated by a central entity deploying third-party LLMAs for Bayesian inference.
@@
There are already many third-party services that offer LLM agents as a service [66], and such agents can be used for the task of performing Bayesian inference on an online platform. Each of these agents has a cost of processing a query associated with it, and each agent can be asked to give more accurate responses (by performing more computations or by using a larger model). Therefore, we can incentivize the LLMAs to share more accurate private information (low-dimensional outputs). However as shown in Section V, LLMAs are still prone to herding, therefore we propose an optimal stopping time formulation for optimally herding and at the same time minimize the cost incurred by the central entity.
IX-B Optimization problem of the central controller with autonomous incentivized LLMAs
We now consider a case where the LLMA are autonomous and are incentivized by a central controller, and the cost function is now dependent on the incentive $p$ as well. Specifically, we consider a cost function $c$ as,

### v2

@@ V-F Summary
Given the privacy, computational, and cost constraints, multiple LLMAs need to interact with each other to perform sequential Bayesian inference on online platforms. This section studied Bayesian social learning in a sequence of LLMAs to analyze multiple LLMAs performing sequential state estimation on online platforms. We discuss the relaxation of the standard Bayesian social learning protocol, in which the agents are allowed to share their private observations. We illustrate the effect on the herding threshold of the public belief when LLMAs are allowed to share private observations and use more accurate LLMs as sensors.
@@ VI Asymmetric Information Structures of Large Language Model Agents
Motivated by the observed model collapse while training LLMs in a sequential manner and data incest in asynchronous Bayesian sensors, this section studies word-of-mouth and asynchronous social learning. Word-of-mouth social learning is a hierarchical social learning paradigm characterized by asymmetric information flow, where lower-level agents process and communicate observations, and top-down influence, where top-level agents dictate public belief. We provide protocols for both kinds of social learning and a corollary that shows that information cascades happen in word-of-mouth learning as well. Techniques to perform stochastic control to prevent model collapse and data incest are not in the scope of this paper and can be considered in future work.
@@
Several recent studies [15] have shown that when LLMs are repeatedly trained on data generated by other LLMs, a phenomenon known as ”model collapse” can occur. In model collapse, the output probability distribution of the model collapses to a degenerate distribution as the model is trained iteratively on data generated by the previously trained model [81]. Similar results have been shown on model distillation, where a smaller LLM is trained using data generated from a larger LLM [82]. We use the Bayesian social learning model to show how an information cascade in LLMAs is similar to the model collapse observed while training LLMs.
@@
Asynchronous social learning in LLMAs is motivated by real-time settings like online platforms where there is a stream of data of the order of a hundred thousand every second [34, 84, 85]. Since LLM functionality within an LLMA often requires several milliseconds to a few seconds, especially if these are third-party services, sequential Bayesian learning is often not possible. This is true particularly when the LLMAs are used for tasks that are more sophisticated than just Bayesian inference[26].
@@ VI-B Word of Mouth Bayesian Social Learning in LLM Agents
We now describe the word-of-mouth social learning protocols in $L$ LLMAs. The protocol is summarized as a pseudo-code in Algorithm 5. The protocol can be considered to run on two timescales. On the slower time scale, the first LLMA receives a new text observation $z_{kL}\sim\mathds{P}(z|x)$ of the state $x$, where $\mathds{P}(z|x)$ is the observation likelihood of text $z$ given the state $x$. The LLMA does not explicitly have knowledge of $\mathds{P}(z|x)$ but only receives a text observation $z$ from it. On the faster time scale, the LLMAs communicate with each other in a sequential fashion by generating text observations corresponding to the low-dimensional features from the received text observation. That is, agent $m$ first receives a text observation $z_{m}$ and uses an LLM to obtain a low dimensional observation $y_{m}$. The agent $m$ then takes an action to update the public belief using (14). Then, the LLMA uses the LLM to generate a synthetic text observation $z_{m+1}$ used by the next agent.
@@
2:     LLMA-($kL$) receives a observation $z_{kL}\sim\mathds{P}(z|x)$.
3:     for $m\in kL,kL+1,kL+2,\dots$$(k+1)L-1$ do
@@
The above result can be proved using Theorem 2, where each agent has a different observation likelihood based on the previous agent, each of which is a concatenation of two observation likelihoods: the LLM as a low-dimensional sensor map and the LLM sampling a text observation from the text observation. Corollary 1 shows that even in the modified protocol of Algorithm 5, where synthetic data is used to aid the decision-making of LLMAs, cascades are inevitable.
@@
Asynchronous Social Learning in LLM Agents: We finally consider the asynchronous social learning setting in LLMAs. Here, the main difference between the previous two topologies is that the agents do not necessarily act in a predefined sequential manner and do not necessarily coordinate.
The protocol is summarized as a pseudo-code in Algorithm 6. The public belief $\pi_{k}$ is updated asynchronously when a LLMA receives a new private observation $z_{k}$, which it parses using the LLM to obtain a low-dimensional observation $y_{k}$. This observation is used to compute the posterior at time $\pi_{k}$ using a previous prior $\pi_{k-1}$. The posterior is then broadcasted.
@@ VII Summary
Part III, therefore, looks at stochastic control for LLMAs. For both regimes of LLMAs, when they are collaborative and autonomous, the paper formulates optimal stopping time problems to control herding by balancing the tradeoff between privacy and estimation. Structural assumptions on the optimal policy of the stopping time problems are proved by making structural assumptions on the cost and observation probabilities. The proposed solutions are extensions to our work in quickest change detection and quickest time herding [4, 86, 16]. A policy gradient algorithm is proposed to estimate the optimal policy without the knowledge of the transition probabilities.
@@ VIII Optimal Stopping Time Control in Centrally Controlled LLMAs
When LLMAs are deployed in real-life settings, they often exhibit bias in their actions, especially when there are multiple such agents [67]. The previous section showed that such bias could be explained by the herding behavior of LLMAs. This section formulates an optimal stopping time problem for quickest time herding in a sequence of LLMAs, which ensures that the herding is optimally delayed by letting the LLMAs share private observations. The delay in herding helps improve the state estimate. We discuss a stochastic control approach to solve the optimal stopping time problem and state assumptions that ensure that the optimal policy for the stopping time problem has a switching threshold curve with respect to the public belief. This structural result is exploited in Section X to efficiently approximate the optimal policy. Finally, we discuss extensions of the problem framework to the optimal switching between models of different sizes in sequential state estimation tasks is discussed. The schematic of the system model considered is illustrated in Figure 10.
@@
VIII-B Social Welfare Objective For Optimal Stopping
@@
This subsection formulates an optimal stopping time problem to delay herding by making the agents optimally switch between two modes, acting benevolently by sharing their private observations or herding by performing the action from (14). Let $a_{k}\in\{0=\texttt{share private observation},1=\texttt{herd}\}$ denote the decision at $k$ for the chosen mode. Let $\mu:\mathcal{P}(\mathcal{X})\to\{0=\texttt{share private observation},1=\texttt{herd}\}$ denotes the stationary policy which maps the public belief to the decision rule. $\mu$ is a sufficient statistic for optimally delaying herding since the information cascade depends on only the public belief [5].For this section, we assume that $\mathcal{X}=\mathcal{U}=\mathcal{Y}$, that is, in the sequential detection task, the observation $y$ is the noisy observation of the underlying state $x$ and $u$ is the maximum a-priori estimate of the state.
@@
Further, we formulate the stopping time problem such that one of the states is of special interest (denoted by $e_{1}$), and a decision to herd is taken once this state is estimated. To formulate the objective for the optimal stopping time problem, we consider the following natural filtration, $\mathcal{F}_{k}$ of the actions and decisions till time $k$, $\mathcal{F}_{k}=\sigma(\{u_{1},\dots,u_{k-1},a_{1},\dots,a_{k}\})$.
@@
The optimal stopping time problem is to decide when to stop sharing private observations ($a_{k}=0$) and announce state $e_{1}$ ($a_{k}=1$). Let $\tau$ denote the stopping time with respect to the filtrations $\mathcal{F}_{k},k=1,\dots$.
We now state the social welfare objective that each LLMA optimizes to solve the optimal stopping time problem and achieve quickest time herding,
$\displaystyle J(\mu)=\mathds{E}_{\mu}$$\displaystyle\left\{\sum_{k=1}^{\tau-1}\rho^{k-1}\mathds{E}\left\{c(x,u_{k})|\mathcal{F}_{k-1}\right\}\right.$ (19)
$\displaystyle\left.+\sum_{k=1}^{\tau-1}\rho^{k-1}d\mathds{E}_{\mu}\left\{\mathds{1}(x=e_{1})|\mathcal{F}_{k-1}\right\}\right.$
$\displaystyle\left.+\ \rho^{\tau-1}\Upsilon\mathds{E}_{\mu}\left\{\mathds{1}(x\neq e_{1})\right\}\right.$
$\displaystyle\left.+\ \frac{\rho^{\tau-1}}{1-\rho}\min_{u\in\mathcal{U}}\mathds{E}\left\{c(x,u)|\mathcal{F}_{\tau-1}\right\}\right\}.$
Here $\rho\in(0,1)$ is the economic discount factor and can be set to a lower value if the central controller only wants the first few actions to determine the policy.
@@
Justification for social welfare cost: In (19), the first part corresponds to the discounted cost incurred if the first $\tau$ agents perform sensing, and the last term corresponds to the cost incurred for the agents after the stopping time $\tau$ (which herd and take the same opportunistic action). The second term is the delay cost with delay parameter $d$ in announcing if the underlying state is $e_{1}$ and the third term is the error cost for misclassification with $\Upsilon$ as the parameter. The social welfare cost incentivizes revealing private observations to delay herding so that enough private observations are available for estimation of the state. The reason we consider estimation with respect to a single state is that many practical applications focus on the identification of a critical state ( hateful user, bad product, etc.). The same social welfare cost function of (19) is optimized to obtain a policy with respect to the belief which decides when to stop herding.
Remark 19.
In a centrally controlled setting, the central controller is the human operator who manages the platform the LLMAs are deployed on. The action space of the agents consists of either performing the optimal action or revealing their private observation. The decision to ‘stop’ corresponds to taking the optimal action, as this would result in herding. Consequently, the optimal stopping formulation inherently leads to agents sharing their private observations until the decision to stop is made.
@@
We consider the decision rule such that depending on the prior, the learners either share the private observation $y_{k}$ or they output the action which minimizes the expected cost of (14) under the public belief, i.e., the action $u_{k}$ is of the form,
@@
The motivation for this decision rule is two-fold: a) the herding can be delayed by ensuring enough private observations are shared, and b) the agents can reduce cost by outputting the opportunistic action once the public belief is strong enough.
@@ VIII-C Structural Results on Optimal Policy
We now state the main structural result on the threshold structure of the optimal policy (21).
@@
Consider the sequential decision problem of LLMAs for detecting state $e_{1}$ with the social welfare cost comprising (19) and the constrained decision rule of (20). Then, under Assumption (S1-S4), the constrained decision rule of (20) is a threshold in the belief space with respect to a threshold switching curve that partitions the belief space $\mathcal{P}(X)$. The optimal policy can be given by,
@@
The above theorem proves that under certain conditions on the cost function and observation matrix (S1-4), the optimal policy for solving the discounted social welfare cost optimal stopping time problem has a switching threshold curve. Such a switching threshold can be approximated by set of lines which can be searched for efficiently using a policy gradient based approach [5]. $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are not known to the LLMA and to the central controller, and the structural result only shows the existence of such disjoint regions. These regions are unknown, and we show in Section X that for $|\mathcal{X}|=2$, one can efficiently search for these regions using a policy gradient algorithm. Although standard reinforcement learning algorithms can be used, we study the extreme case to derive an efficient algorithm. This is computationally tractable, unlike optimizing for a non-structured policy where a finite approximation is used for a general infinite-dimensional policy. This is computationally tractable, in contrast to optimizing for a non-structured policy since it’s a finite approximation of a general infinite-dimensional policy.
Remark 20.
The LLM agents share their current private observations, and the stopping-time policy determines when the sequence of agents ceases sharing private information. This approach can delay herding while ensuring that not all private information is disclosed. The privacy loss can be quantified using an information-theoretic metric that captures the exact loss of private information. Investigating this in practical scenarios would be an intriguing direction for future research.
@@ VIII-D Optimal Switching Between Different LLMs
The formulation that we proposed in this section can also be used by the LLMAs to switch between LLMs of different sizes. LLMs with a higher number of parameters can process more tokens and give more accurate responses; however, they are more expensive and use more computational resources. Therefore we can use an action space $\mathcal{U}=\{1=\mathcal{L}_{\text{small}},\mathcal{L}_{\text{large}}\}$, where $\mathcal{L}_{\text{large}}$ is an LLM with significantly (1010 1010$\times$) more parameters than $\mathcal{L}_{\text{small}}$.
@@ VIII-E Summary
When a central entity is designing and deploying LLMAs for sequential Bayesian inference on online platforms, the entity needs to ensure such LLMAs do not have a bias due to herding. In this section, we solved the problem of LLMAs optimally herding in Bayesian social learning framework to announce a particular state, such that their opportunistic cost and cost of obtaining and sharing the observation is balanced. The Bayesian agents considered in this section were cooperative and shared the same socialistically optimal policy. We also briefly discussed how such a scheme can be used to optimally switch between different sizes of LLMs to achieve an optimal tradeoff between cost and accuracy of estimation. In the next section we consider the problem in a different setting, where a central controller can incentivize autonomous LLMAs, and needs to control the incentives for improved state estimation.
@@ IX Optimal Stopping Time Control for Autonomous LLMAs
The LLMAs used for Bayesian inference can often be from different third-party services, each of which requires an incentive to perform the task. Motivated by controlling bias in such LLMAs, this section considers the problem of the central learner optimally incentivizing a sequence of autonomous LLMAs to delay their herding and obtain more accurate estimates. We first formulate the optimization problem of the central controller as a discounted cost POMDP and then show that under structural assumptions on the cost and observation matrix, the optimal policy has a threshold structure with respect to the public belief. This structure is exploited in Algorithm 7 to approximate the optimal incentivization policy of the central controller. The schematic of the setup considered in this section is illustrated in Figure 11.
@@
Figure 11: Schematic of stochastic control approach for optimally delaying herding in a sequence of incentivized autonomous LLMAs. This is motivated by a central entity deploying third-party LLMAs for Bayesian inference.
@@
There are already several third-party services that offer LLM agents as a service [17], and such agents can be used for the task of performing Bayesian inference on an online platform. Each of these agents has a cost of processing a query associated with it, and each agent can be asked to give more accurate responses (by performing more computations or by using a larger model). Therefore, we can incentivize the LLMAs to share more accurate private information (low-dimensional outputs). However as shown in Section V, LLMAs are still prone to herding, therefore we propose an optimal stopping time formulation for optimally herding and at the same time minimize the cost incurred by the central entity.
IX-B Optimization problem of the central controller with incentivized LLMAs
We now consider a case where the LLMA are incentivized by a central controller, and the cost function is now dependent on the incentive $p$ as well. Specifically, we consider a cost function $c$ as,